Use appropriate algebra and Theorem 7.2.1 to find the given inverse Laplace transform. (Write your answer as a function of t.)

ℒ−1{(s/(s2 + 3s − 4)}


1
Expert's answer
2021-04-15T06:56:35-0400

"\frac{s}{s^2+3s-4}=\frac{s}{(s-1)(s+4)}="


"=\frac{A}{s-a}+\frac{B}{s+4}=\frac{(A+B)s+4A-B}{(s-1)(s+4)}"


we have a system:


"\begin{cases}
A+B=1\\
4A-B=0
\end{cases}"


"5A=1"

"A=\frac{1}{5}"

"B=\frac{4}{5}"


"\frac{s}{s^2+3s-4}=\frac{1}{5(s-1)}+\frac{4}{5(s+4)}"


"L^{-1}(\frac{s}{s^2+3s-4})=L^{-1}(\frac{1}{5(s-1)})+L^{-1}(\frac{4}{5(s+4)})="


"=\frac{1}{5}e^t+\frac{4}{5}e^{-4t}"